### 19 Mar 2026 v2

# The pattern block method for generating random numbers : Reformulation and generalization of the Ziggurat method using conditional random variables

Abstract
The Ziggurat method is an efficient rejection sampling method for generating 11-dimensional normally distributed random numbers. This paper proposes the pattern block method, which is a generalization of the Ziggurat method. The pattern block method can be used to generate random numbers from multimodal density functions or multidimensional distributions. We demonstrate the effectiveness of the pattern block method through several examples.
Keywords: Random number generation, The Ziggurat method, The pattern block method. 0002020 Mathematics Subject Classification: Primary 65C10; Secondary 62-08.
1 Introduction
Random number generation is a foundational component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness through numerical examples. However, the random variable underlying the Ziggurat algorithm has not been defined. Therefore, in this study, we reformulate and generalize the Ziggurat method using conditional random variables. Conditional random variables have played an important role in [7, 8, 9, 10].
This paper is organized as follows. In Section 2, we present the pattern block method proposed in this study. Sections 3 and 4 present applications of the pattern block method. Section 5 concludes.
2 The pattern block method
Let $(E,\mathcal{E},\mu)$ be a $\sigma$-finite measure space. Let $f:E\rightarrow\mathbb{R}$ be an $\mathcal{E}/\mathcal{B}(\mathbb{R})$-measurable function that satisfies
$f(x)\geq 0\ (x\in E),\qquad K:=\int_{E}f(x)\,\mu(dx)\in(0,\infty),$ (1)
where $\mathcal{B}(\mathbb{R})$ is the Borel $\sigma$-algebra on $\mathbb{R}$. Then $\frac{1}{K}f(x)$ is a probability density function. Let
$\Gamma[f]:=\left\{(x,y)\in E\times\mathbb{R}\ \big|\ 0\leq y\leq f(x)\right\}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R}),$
where $\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ denotes the product $\sigma$- algebra of $\mathcal{E}$ and $\mathcal{B}(\mathbb{R})$. Let $\nu$ be the measure on $(E\times\mathbb{R},\mathcal{E}\otimes\mathcal{B}(\mathbb{R}))$
$\nu(C):=\iint_{E\times\mathbb{R}}1_{C}(x,y)\,\mu(dx)\,dy\ \,(C\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})).$ (2)
Let $\widetilde{\pi}:E\times\mathbb{R}\to E$ be the projection defined by
$\widetilde{\pi}(x,y):=x\quad((x,y)\in E\times\mathbb{R}).$
Let $N\in\mathbb{N}$. Assume that $B_{1},B_{2},\cdots,B_{N}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ and $B:=\bigcup_{i=1}^{N}B_{i}$ satisfy the following three conditions:
$\nu(B_{i})>0\quad(i=1,2,\cdots,N),$ (3)
$\nu(B_{i}\cap B_{j})=0\quad(i\neq j),$ (4)
$\nu\left(\Gamma[f]\setminus B\right)=0,\quad\nu(B)<\infty.$ (5)
In this study, we call $B_{1},B_{2},\cdots,B_{N}$ “pattern block of f(x)f(x).” Using (1), (4), (5), we obtain
$K=\nu(\Gamma[f])\leq\nu(B)=\sum_{i=1}^{N}\nu(B_{i})<\infty.$ (6)
Figure 1: $f(x)$ and its pattern block
Let $\Psi$ be the block selection function defined by
$\Psi(u):=\min\left\{i\in\{1,2,\cdots,N\}\ \middle|\ u<\sum_{j=1}^{i}\frac{\nu(B_{j})}{\nu(B)}\right\}\qquad(0\leq u<1).$
For $i=1,2,\cdots,N$, let
$B_{i}[f]:=B_{i}\cap\Gamma[f]=\left\{(x,y)\in B_{i}\ |\ 0\leq y\leq f(x)\right\}.$
Let $\Lambda$ be the acceptance set defined by
$\Lambda:=\big\{(u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N})\in[0,1)\times E^{N}\times\mathbb{R}^{N}\ \big|\ (v_{\Psi(u)},w_{\Psi(u)})\in B_{\Psi(u)}[f]\big\}.$ (7)
Let $\pi_{\Psi}:[0,1)\times E^{N}\times\mathbb{R}^{N}\to E$ be the projection
$\pi_{\Psi}(u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N}):=\widetilde{\pi}(v_{\Psi(u)},w_{\Psi(u)})=v_{\Psi(u)}$ (8)
$((u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N})\in[0,1)\times E^{N}\times\mathbb{R}^{N}).$
Let $(\Omega,\mathcal{F},P)$ be a probability space. Let $U,W_{1},W_{2},\cdots,W_{N}$ be 11-dimensional random variables on $(\Omega,\mathcal{F},P)$, and let $V_{1},V_{2},\cdots,V_{N}$ be $(E,\mathcal{E})$-valued random variables on $(\Omega,\mathcal{F},P)$. For $i=1,2,\cdots,N$, we assume that $(V_{i},W_{i})$ and $U$ are independent and satisfy
$P((V_{i},W_{i})\in C)=\dfrac{\nu(C\cap B_{i})}{\nu(B_{i})}\quad(C\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})),$ (9)
$P\left(U\in D\right)=\int_{D}1_{[0,1)}(u)\,du\qquad(D\in\mathcal{B}(\mathbb{R})).$ (10)
Let $Z:(\Omega,\mathcal{F},P)\rightarrow(\mathbb{R}\times E^{N}\times\mathbb{R}^{N},\ \mathcal{B}(\mathbb{R})\otimes\mathcal{E}\otimes\mathcal{E}\otimes\cdots\otimes\mathcal{E}\otimes\mathcal{B}(\mathbb{R}^{N}))$ be the random variable given by
$Z:=(U,V_{1},V_{2},\cdots,V_{N},W_{1},W_{2},\cdots,W_{N}).$ (11)
Then we obtain the following lemma.
Lemma 2.1.
For $A\in\mathcal{E}$, we have
$P(V_{\Psi(U)}\in A,Z\in\Lambda)=\frac{1}{\nu(B)}\int_{A}f(x)\ \mu(dx).$ (12)
Proof.
Since $(V_{i},W_{i})$ and $U$ are independent for every $i=1,2,\cdots,N$, we obtain
$P(V_{\Psi(U)}\in A,Z\in\Lambda)=\sum_{i=1}^{N}P(\Psi(U)=i)\cdot P((V_{i},W_{i})\in B_{i}[f]\cap(A\times\mathbb{R})).$ (13)
It holds from (4), (9), (10) and (13) that
$P(V_{\Psi(U)}\in A,Z\in\Lambda)$ $=\sum_{i=1}^{N}\frac{\nu(B_{i})}{\nu(B)}\cdot E[1_{B_{i}[f]\cap(A\times\mathbb{R})}(V_{i},W_{i})]$
$=\sum_{i=1}^{N}\frac{\nu(B_{i})}{\nu(B)}\cdot\frac{1}{\nu(B_{i})}\iint_{B_{i}\cap(A\times\mathbb{R})}1_{[0,f(x)]}(y)\,\mu(dx)\,dy$
$=\frac{1}{\nu(B)}\iint_{B\cap(A\times\mathbb{R})}1_{[0,f(x)]}(y)\,\mu(dx)\,dy.$ (14)
Using (5), (14) and the Fubini’s theorem, we obtain (12). ∎
By Lemma 2.1, (1) and (6), it holds that
$P(Z\in\Lambda)=\dfrac{1}{\nu(B)}\displaystyle\int_{E}f(x)\ \mu(dx)=\dfrac{K}{\nu(B)}>0.$ (15)
Let $P_{Z^{-1}(\Lambda)}$ be the probability measure on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F})$ defined by
$P_{Z^{-1}(\Lambda)}(A):=\frac{P(A)}{P(Z\in\Lambda)},\qquad A\in Z^{-1}(\Lambda)\cap\mathcal{F}:=\{Z^{-1}(\Lambda)\cap F\ |\ F\in\mathcal{F}\}.$
Let $Z|_{\Lambda}$ denote the restriction of $Z$ to $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$
$Z:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(\Lambda,\mathcal{B}(\Lambda)).$
Then $Z|_{\Lambda}$ is a random variable. Let $X:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(E,\mathcal{E})$ be the random variable defined by
$X:=\pi_{\Psi}(Z|_{\Lambda})=\pi_{\Psi}((U,V_{1},V_{2},\cdots,V_{N},W_{1},W_{2},\cdots,W_{N})|_{\Lambda}).$ (16)
$Z|_{\Lambda}$ and $X$ are random variables defined on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$, and are called conditional random variables. Using Lemma 2.1 and (15), we obtain the following theorem.
Theorem 2.1.
For $A\in\mathcal{E}$, we have
$P_{Z^{-1}(\Lambda)}(X\in A)=\frac{1}{K}\int_{A}f(x)\ \mu(dx).$
Thus, using the random variable $\pi_{\Psi}(Z|_{\Lambda})$, we can generate random numbers from the probability density function $\frac{1}{K}f(x)$. We call this technique the pattern block method. The algorithm of the pattern block method is as follows:
  • Step 1.
    Generate $U$ and $(V_{\Psi(U)},W_{\Psi(U)})$.
  • Step 2.
    If $W_{\Psi(U)}\leq f(V_{\Psi(U)})$, return $V_{\Psi(U)}$.
  • Step 3.
    If $W_{\Psi(U)}>f(V_{\Psi(U)})$, go to Step 1.
Let $R_{{\rm adoption}}$ be the adoption rate defined by
$R_{{\rm adoption}}:=\frac{K}{\nu(B_{1})+\nu(B_{2})+\cdots+\nu(B_{N})}.$ (17)
3 Applications to generate 11-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is $(\mathbb{R},\mathcal{B}(\mathbb{R}))$ and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then $\nu$ defined by (2) is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
3.1 The Ziggurat method
Figure 2: Ziggurat method
In this subsection, we explain that the pattern block method is a generalization of the Ziggurat method. Assume that a probability density function $f(x)$ is strictly decreasing for $x>0$ and satisfies $\int_{0}^{\infty}f(x)\ dx=1$. Let $N\in\mathbb{N}$ with $N\geq 2$, and let $0=x_{0}<x_{1}<\cdots<x_{N-1}$. Let
$B_{i}=[0,x_{i}]\times[f(x_{i}),f(x_{i-1})]\qquad(i=1,2,\cdots,N-1),$
$B_{N}=([0,x_{N-1}]\times[0,f(x_{N-1})])\cup\{(x,y)\in\mathbb{R}^{2}\ |\ x\geq x_{N-1},0\leq y\leq f(x)\}$
be the pattern block of $f(x)$. Then the pattern block method corresponds to the Ziggurat method.
3.2 The density function with singularities
Let $\varphi(x)=\dfrac{1}{\pi\sqrt{x(1-x)}}$ and $h(x)=1+\sin(8\pi x)$. Let $f(x)$ be the density function
$f(x)=\left\{\begin{array}[]{cl}h(x)\varphi(x)&(0<x<1),\\ 0&(\text{otherwise}).\end{array}\right$.
Let $N=8$. We define the pattern block of $f(x)$ as
$B_{i}=\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;a_{i-1}\leq x\leq a_{i},\;0\leq y\leq b_{i}\varphi(x)\right\}$
for $i=1,2,\cdots,8$, where
$a_{i}:=\dfrac{i}{8},\quad b_{i}:=\left\{\begin{array}[]{cl}2&(\text{when $i$ is odd})\\ 1&(\text{when $i$ is even})\end{array}\right.(1\leq i\leq 8).$
Fig. 3 shows $f(x)$ and $B_{1},B_{2},\cdots,B_{8}$.
Figure 3: $f(x)$ and its pattern block $B_{1},B_{2},\cdots,B_{8}$
Since $\varphi(x)$ is the density function of the Beta distribution $B\left(\tfrac{1}{2},\tfrac{1}{2}\right)$, its cumulative distribution function $\Phi(x)$ is given by
$\Phi(x)=\frac{2}{\pi}\arcsin\left(\sqrt{x}\right)\qquad(0<x<1).$
Then $F_{V_{i}}(x):=P(V_{i}\leq x)$ is given by
$F_{V_{i}}(x)=\left\{\begin{array}[]{cl}0&(x\leq a_{i-1}),\\ \dfrac{b_{i}}{|B_{i}|}\bigl(\Phi(x)-\Phi(a_{i-1})\bigr)&(a_{i-1}\leq x\leq a_{i}),\\ 1&(x\geq a_{i}).\end{array}\right$.
Assume that the independent random variables $U$, $\xi$, $\eta$ defined on a probability space $(\Omega,\mathcal{F},P)$ are uniformly distributed on $[0,1)$. We define the random variables $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ as
$V_{i}:=$ $F_{V_{i}}^{-1}(\xi)=\Phi^{-1}(\Phi(a_{i-1})+\xi\,\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr))=\sin^{2}\!\left(\frac{\pi}{2}\left\{\Phi(a_{i-1})+\xi\,\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr)\right\}\right),$
$W_{i}:=$ $b_{i}\varphi(V_{i})\,\eta=b_{i}\varphi\big(F_{V_{i}}^{-1}(\xi)\big)\,\eta.$
Then $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ satisfy (9). The adoption rate can be calculated as follows:
$R_{{\rm adoption}}=\frac{1}{\sum_{i=1}^{8}b_{i}\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr)}=\dfrac{2}{3}.$
Thus, by applying the pattern block method, we obtain $10,000$ random numbers from the probability density function $f(x)$. To obtain $10,000$ random numbers, $15,048$ random number triplets $(U,\xi,\eta)$ were required. Fig. 4 shows the density scale histogram of $10,000$ random numbers.
Figure 4: Histogram and density function
4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is given by $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. For every $C\in\mathcal{B}(E\times\mathbb{R})$, $|C|$ denotes the 33-dimensional Lebesgue measure $\nu(C)$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$
$f(x_{1},x_{2})=c\Big\{e^{-x_{1}^{2}-x_{2}^{2}}+\dfrac{1}{2}e^{-(x_{1}-2)^{2}-(x_{2}-2)^{2}}\Big\},$
where $c=\dfrac{2119}{9970}$. Let $b_{0}=\cfrac{1}{40}$, $b_{1}=\dfrac{1}{15}$, $b_{2}=c\big(\dfrac{1}{e^{8}}+\dfrac{1}{2}\big)$, $b_{3}=c\big(1+\dfrac{1}{2e^{8}}\big)$. We define $\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$ as
$\big[\underline{\kappa}_{1},\overline{\kappa}_{1}\big]:=\big[0,b_{0}\big],\quad\big[\underline{\kappa}_{2},\overline{\kappa}_{2}\big]:=\big[b_{0},b_{1}\big],\quad\big[\underline{\kappa}_{3},\overline{\kappa}_{3}\big]:=\big[b_{1},b_{2}\big],\quad\big[\underline{\kappa}_{4},\overline{\kappa}_{4}\big]:=\big[b_{1},b_{2}\big],\quad\big[\underline{\kappa}_{5},\overline{\kappa}_{5}\big]:=\big[b_{2},b_{3}\big].$
Moreover, we define $(c_{i,1},c_{i,2})$, $d_{i}$ $(i=3,4,5)$ as
$(c_{3,1},c_{3,2})=(c_{5,1},c_{5,2})=(0,0),\quad(c_{4,1},c_{4,2})=(2,2),\quad d_{3}=\dfrac{5}{4},\quad d_{4}=d_{5}=1.$
Let $N=5$. We define the pattern block of $f(x_{1},x_{2})$ as $B_{i}=E_{i}\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$, where
$E_{1}=E,\quad E_{2}=\left\{(x_{1},x_{2})\in E\ \middle|\ f(x_{1},x_{2})\geq b_{0}\right\},$
$E_{i}=\left\{(x_{1},x_{2})\in E\ \middle|\ (x_{1}-c_{i,1})^{2}+(x_{2}-c_{i,2})^{2}\leq d_{i}^{2}\right\}\qquad(i=3,4,5).$
Figure 5: $f(x_{1},x_{2})=b_{0}$
Figure 6: $f(x_{1},x_{2})=b_{1}$
Figure 7: $f(x_{1},x_{2})=b_{2}$
Assume that the independent random variables $U$, $\xi_{1}$, $\xi_{2}$, $\eta$ defined on a probability space $(\Omega,\mathcal{F},P)$ are uniformly distributed on $[0,1)$. Let
$V_{1}^{(j)}=\underline{a}+(\overline{a}-\underline{a})\cdot\xi_{j}\quad(j=1,2),\quad W_{1}=b_{0}\cdot\eta.$ (18)
Then $(V_{1}^{(1)},V_{1}^{(2)},W_{1})$ satisfies (9). Let
$\widetilde{V}_{2}^{(1)}:=\dfrac{11}{2}\xi_{1}-2,\quad\widetilde{V}_{2}^{(2)}:=\dfrac{11}{2}\xi_{2}-2.$
We define $(V_{2}^{(1)},V_{2}^{(2)},W_{2})$ as
$(V_{2}^{(1)},V_{2}^{(2)}):=\left(\widetilde{V}_{2}^{(1)},\widetilde{V}_{2}^{(2)}\right)\big|_{E_{2}},\quad W_{2}:=(b_{1}-b_{0})\eta+b_{0}.$ (19)
Then $(V_{2}^{(1)},V_{2}^{(2)},W_{2})$ satisfies (9). Let
$\Phi_{i}(r,\theta,y):=(c_{i,1}+r\cos\theta,c_{i,2}+r\sin\theta,y),\quad\widetilde{B}_{i}:=\left[0,d_{i}\right]\times[0,2\pi)\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]\qquad(i=3,4,5).$
Then we have $B_{i}=\Phi_{i}(\widetilde{B}_{i})$ $(i=3,4,5)$. Let
$F_{R_{i}}(r):=\left\{\begin{array}[]{cl}0&(r\leq 0)\\ \dfrac{\pi r^{2}\big(\overline{\kappa}_{i}-\underline{\kappa}_{i}\big)}{|B_{i}|}&\left(0\leq r\leq d_{i}\right)\\ 1&\left(r\geq d_{i}\right).\end{array}\right$.
We define $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ as
$V_{i}^{(1)}=c_{i,1}+F_{R_{i}}^{-1}(\xi_{1})\cos 2\pi\xi_{2},\quad V_{i}^{(2)}=c_{i,2}+F_{R_{i}}^{-1}(\xi_{1})\sin 2\pi\xi_{2},\quad W_{i}=(\overline{\kappa}_{i}-\underline{\kappa}_{i})\eta+\underline{\kappa}_{i}\quad(i=3,4,5).$ (20)
Then $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ satisfy (9). Thus, by applying the pattern block method, we obtain $500,000$ random vectors from the probability density function $f(x_{1},x_{2})$. To obtain $500,000$ random vectors, $1,375,334$ random number quartets $(U,V_{\Psi(U)}^{(1)},V_{\Psi(U)}^{(2)},W_{\Psi(U)})$ were required. The adoption rate can be calculated as follows:
$R_{\rm{adoption}}=\dfrac{1}{\sum_{i=1}^{5}|B_{i}|}=0.3644\cdots.$
Fig. 9 and Fig. 9 show the density scale histogram of $500,000$ random vectors and the density function $y=f(x_{1},x_{2})$, respectively.
Figure 8: Histogram
Figure 9: $y=f(x_{1},x_{2})$
5 Conclusion
We introduced the pattern block method, which is a generalization of the Ziggurat method. In our future work, we aim to develop an algorithm that identifies pattern block of a given density function with high adoption rate and low search cost.
Acknowledgments
The authors would like to thank Prof. Masaaki Fukasawa (Osaka University) and Prof. Suguru Yamanaka (Aoyama Gakuin University) for their helpful comments and discussions on the subject matter. This work was supported by JSPS KAKENHI Grant Numbers JP22K01556 and JP25K05168.