### 24 Mar 2026 v3

# Pattern block method for generating random numbers : Reformulation and generalization of the Ziggurat method using conditional random variables

Abstract
The Ziggurat method is an efficient rejection sampling technique for generating one-dimensional normally distributed random numbers. This study proposes the pattern block method, a generalization of the Ziggurat method. The pattern block method enables the generation of random numbers from multimodal density functions and multidimensional distributions. The effectiveness of the pattern block method is demonstrated through several examples.
Keywords: Random number generation, Ziggurat method, Pattern block method. 0002020 Mathematics Subject Classification: Primary 65C10; Secondary 62-08.
1 Introduction
Random number generation is a fundamental component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness using numerical examples. However, the random variable underlying the Ziggurat algorithm has not yet been formally defined. Therefore, this study reformulates and generalizes the Ziggurat method using conditional random variables. Conditional random variables play an important role in [7, 8, 9, 10].
The remainder of this paper is organized as follows. Section 2 presents the pattern block method proposed in this study. Sections 3 and 4 present applications of the pattern block method. Finally, Section 5 concludes.
2 The pattern block method
Let $(E,\mathcal{E},\mu)$ be a $\sigma$-finite measure space. Let $f:E\rightarrow\mathbb{R}$ be a $\mathcal{E}/\mathcal{B}(\mathbb{R})$-measurable function that satisfies
$f(x)\geq 0\ (x\in E),\qquad K:=\int_{E}f(x)\,\mu(dx)\in(0,\infty),$ (1)
where $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra of $\mathbb{R}$. Then $\frac{1}{K}f(x)$ is a probability density function. Let
$\Gamma[f]:=\left\{(x,y)\in E\times\mathbb{R}\ \big|\ 0\leq y\leq f(x)\right\}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R}),$
where $\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ denotes the product $\sigma$- algebra of $\mathcal{E}$ and $\mathcal{B}(\mathbb{R})$. Let $\nu$ be the measure on $(E\times\mathbb{R},\mathcal{E}\otimes\mathcal{B}(\mathbb{R}))$
$\nu(C):=\iint_{E\times\mathbb{R}}1_{C}(x,y)\,\mu(dx)\,dy\ \,(C\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})).$ (2)
Let $\widetilde{\pi}:E\times\mathbb{R}\to E$ be a projection defined by
$\widetilde{\pi}(x,y):=x\quad((x,y)\in E\times\mathbb{R}).$
Let $N\in\mathbb{N}$. We assume that $B_{1},B_{2},\cdots,B_{N}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ and $B:=\bigcup_{i=1}^{N}B_{i}$ satisfy the following three conditions:
$\nu(B_{i})>0\quad(i=1,2,\cdots,N),$ (3)
$\nu(B_{i}\cap B_{j})=0\quad(i\neq j),$ (4)
$\nu\left(\Gamma[f]\setminus B\right)=0,\quad\nu(B)<\infty.$ (5)
We refer to $B_{1},B_{2},\cdots,B_{N}$ as the pattern blocks of $f(x)$. Using (1), (4), and (5), we obtain
$K=\nu(\Gamma[f])\leq\nu(B)=\sum_{i=1}^{N}\nu(B_{i})<\infty.$ (6)
Figure 1: $f(x)$ and its pattern blocks
Let $\Psi$ be the block selection function defined as
$\Psi(u):=\min\left\{i\in\{1,2,\cdots,N\}\ \middle|\ u<\sum_{j=1}^{i}\frac{\nu(B_{j})}{\nu(B)}\right\}\qquad(0\leq u<1).$
For $i=1,2,\cdots,N$, let
$B_{i}[f]:=B_{i}\cap\Gamma[f]=\left\{(x,y)\in B_{i}\ |\ 0\leq y\leq f(x)\right\}.$
Let $\Lambda$ be the acceptance set defined as
$\Lambda:=\big\{(u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N})\in[0,1)\times E^{N}\times\mathbb{R}^{N}\ \big|\ (v_{\Psi(u)},w_{\Psi(u)})\in B_{\Psi(u)}[f]\big\}.$ (7)
Let $\pi_{\Psi}:[0,1)\times E^{N}\times\mathbb{R}^{N}\to E$ denote the projection:
$\pi_{\Psi}(u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N}):=\widetilde{\pi}(v_{\Psi(u)},w_{\Psi(u)})=v_{\Psi(u)}$ (8)
$((u,v_{1},v_{2},\cdots,v_{N},w_{1},w_{2},\cdots,w_{N})\in[0,1)\times E^{N}\times\mathbb{R}^{N}).$
Let $(\Omega,\mathcal{F},P)$ be a probability space. Let $U,W_{1},W_{2},\cdots,W_{N}$ be 11-dimensional random variables on $(\Omega,\mathcal{F},P)$, and let $V_{1},V_{2},\cdots,V_{N}$ be $(E,\mathcal{E})$-valued random variables on $(\Omega,\mathcal{F},P)$. For $i=1,2,\cdots,N$, we assume that $(V_{i},W_{i})$ and $U$ are independent and satisfy
$P((V_{i},W_{i})\in C)=\dfrac{\nu(C\cap B_{i})}{\nu(B_{i})}\quad(C\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})),$ (9)
$P\left(U\in D\right)=\int_{D}1_{[0,1)}(u)\,du\qquad(D\in\mathcal{B}(\mathbb{R})).$ (10)
Let $Z:(\Omega,\mathcal{F},P)\rightarrow(\mathbb{R}\times E^{N}\times\mathbb{R}^{N},\ \mathcal{B}(\mathbb{R})\otimes\mathcal{E}^{\otimes N}\otimes\mathcal{B}(\mathbb{R}^{N}))$ be a random variable given by
$Z:=(U,V_{1},V_{2},\cdots,V_{N},W_{1},W_{2},\cdots,W_{N}).$ (11)
Then we obtain the following lemma.
Lemma 2.1.
For $A\in\mathcal{E}$, we have
$P(V_{\Psi(U)}\in A,Z\in\Lambda)=\frac{1}{\nu(B)}\int_{A}f(x)\ \mu(dx).$ (12)
Proof.
Since $(V_{i},W_{i})$ and $U$ are independent for every $i=1,2,\cdots,N$, we obtain
$P(V_{\Psi(U)}\in A,Z\in\Lambda)=\sum_{i=1}^{N}P(\Psi(U)=i)\cdot P((V_{i},W_{i})\in B_{i}[f]\cap(A\times\mathbb{R})).$ (13)
From (4), (9), (10), and (13), we obtain
$P(V_{\Psi(U)}\in A,Z\in\Lambda)$ $=\sum_{i=1}^{N}\frac{\nu(B_{i})}{\nu(B)}\cdot E[1_{B_{i}[f]\cap(A\times\mathbb{R})}(V_{i},W_{i})]$
$=\sum_{i=1}^{N}\frac{\nu(B_{i})}{\nu(B)}\cdot\frac{1}{\nu(B_{i})}\iint_{B_{i}\cap(A\times\mathbb{R})}1_{[0,f(x)]}(y)\,\mu(dx)\,dy$
$=\frac{1}{\nu(B)}\iint_{B\cap(A\times\mathbb{R})}1_{[0,f(x)]}(y)\,\mu(dx)\,dy.$ (14)
Using (5), (14), and Fubini’s theorem, we obtain (12). ∎
By Lemma 2.1, (1), and (6), it follows that
$P(Z\in\Lambda)=\dfrac{1}{\nu(B)}\displaystyle\int_{E}f(x)\ \mu(dx)=\dfrac{K}{\nu(B)}>0.$ (15)
Let $P_{Z^{-1}(\Lambda)}$ be the probability measure on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F})$ defined by
$P_{Z^{-1}(\Lambda)}(A):=\frac{P(A)}{P(Z\in\Lambda)},\qquad A\in Z^{-1}(\Lambda)\cap\mathcal{F}:=\{Z^{-1}(\Lambda)\cap F\ |\ F\in\mathcal{F}\}.$
Let $Z|_{\Lambda}$ denote the restriction of $Z$ to $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$
$Z:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(\Lambda,\mathcal{B}(\Lambda)).$
Thus, $Z|_{\Lambda}$ is a random variable. Let $X:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(E,\mathcal{E})$ be a random variable defined as follows:
$X:=\pi_{\Psi}(Z|_{\Lambda})=\pi_{\Psi}((U,V_{1},V_{2},\cdots,V_{N},W_{1},W_{2},\cdots,W_{N})|_{\Lambda}).$ (16)
$Z|_{\Lambda}$ and $X$ are conditional random variables defined on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$. Using Lemma 2.1 and (15), we obtain the following theorem:
Theorem 2.1.
For $A\in\mathcal{E}$, we have
$P_{Z^{-1}(\Lambda)}(X\in A)=\frac{1}{K}\int_{A}f(x)\ \mu(dx).$
Thus, using the conditional random variable $\pi_{\Psi}(Z|_{\Lambda})$, we can generate random numbers from the density function $\frac{1}{K}f(x)$. We call this technique the pattern block method. The algorithm is as follows:
  • Step 1.
    Generate $U$ and $(V_{\Psi(U)},W_{\Psi(U)})$.
  • Step 2.
    If $W_{\Psi(U)}\leq f(V_{\Psi(U)})$, return $V_{\Psi(U)}$.
  • Step 3.
    If $W_{\Psi(U)}>f(V_{\Psi(U)})$, go to Step 1.
Let $R_{{\rm adoption}}$ be the adoption rate defined as
$R_{{\rm adoption}}:=\frac{K}{\nu(B_{1})+\nu(B_{2})+\cdots+\nu(B_{N})}.$ (17)
3 Applications to generate 11-dimensional random numbers
In this section, we assume the following. $(E,\mathcal{E})$ is $(\mathbb{R},\mathcal{B}(\mathbb{R}))$, and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then, $\nu$, defined by (2), is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
3.1 The Ziggurat method
Figure 2: Ziggurat method
In this subsection, we explain that the pattern block method is a generalization of the Ziggurat method. We assume that the probability density function $f(x)$ strictly decreases for $x>0$ and satisfies $\int_{0}^{\infty}f(x)\ dx=1$. Let $N\in\mathbb{N}$ with $N\geq 2$, and let $0=x_{0}<x_{1}<\cdots<x_{N-1}$. Let
$B_{i}=[0,x_{i}]\times[f(x_{i}),f(x_{i-1})]\qquad(i=1,2,\cdots,N-1),$
$B_{N}=([0,x_{N-1}]\times[0,f(x_{N-1})])\cup\{(x,y)\in\mathbb{R}^{2}\ |\ x\geq x_{N-1},0\leq y\leq f(x)\}$
be the pattern blocks of $f(x)$. The pattern block method then corresponds to the Ziggurat method.
3.2 Density function with singularities
Let $\varphi(x)=\dfrac{1}{\pi\sqrt{x(1-x)}}$ and $h(x)=1+\sin(8\pi x)$. Let $f(x)$ be the density function defined as
$f(x)=\left\{\begin{array}[]{cl}h(x)\varphi(x)&(0<x<1),\\ 0&(\text{otherwise}).\end{array}\right$.
Let $N=8$. The pattern blocks of $f(x)$ are defined as
$B_{i}=\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;a_{i-1}\leq x\leq a_{i},\;0\leq y\leq b_{i}\varphi(x)\right\}$
for $i=1,2,\cdots,8$, where
$a_{i}:=\dfrac{i}{8},\quad b_{i}:=\left\{\begin{array}[]{cl}2&(\text{when $i$ is odd})\\ 1&(\text{when $i$ is even})\end{array}\right.(1\leq i\leq 8).$
Fig. 3 shows $f(x)$ and $B_{1},B_{2},\cdots,B_{8}$.
Figure 3: $f(x)$ and its pattern blocks $B_{1},B_{2},\cdots,B_{8}$
Because $\varphi(x)$ is the density function of the Beta distribution $B\left(\tfrac{1}{2},\tfrac{1}{2}\right)$, its cumulative distribution function $\Phi(x)$ is given by
$\Phi(x)=\frac{2}{\pi}\arcsin\left(\sqrt{x}\right)\qquad(0<x<1).$
Then $F_{V_{i}}(x):=P(V_{i}\leq x)$ is given by
$F_{V_{i}}(x)=\left\{\begin{array}[]{cl}0&(x\leq a_{i-1}),\\ \dfrac{b_{i}}{|B_{i}|}\bigl(\Phi(x)-\Phi(a_{i-1})\bigr)&(a_{i-1}\leq x\leq a_{i}),\\ 1&(x\geq a_{i}).\end{array}\right$.
We assume that the independent random variables $U$, $\xi$, and $\eta$, defined on the probability space $(\Omega,\mathcal{F},P)$, are uniformly distributed on $[0,1)$. We define the random variables $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ as
$V_{i}:=$ $F_{V_{i}}^{-1}(\xi)=\Phi^{-1}(\Phi(a_{i-1})+\xi\,\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr))=\sin^{2}\!\left(\frac{\pi}{2}\left\{\Phi(a_{i-1})+\xi\,\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr)\right\}\right),$
$W_{i}:=$ $b_{i}\varphi(V_{i})\,\eta=b_{i}\varphi\big(F_{V_{i}}^{-1}(\xi)\big)\,\eta.$
Subsequently, $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ satisfy (9). The adoption rate is calculated as follows:
$R_{{\rm adoption}}=\frac{1}{\sum_{i=1}^{8}b_{i}\bigl(\Phi(a_{i})-\Phi(a_{i-1})\bigr)}=\dfrac{2}{3}.$
Thus, by applying the pattern block method, we obtained $10000$ random numbers from the probability density function $f(x)$. To obtain $10000$ random numbers, $15048$ random number triplets $(U,\xi,\eta)$ were required. Fig. 4 shows the density scale histogram for $10000$ random numbers.
Figure 4: Histogram and density function
4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is defined as follows: $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. For every $C\in\mathcal{B}(E\times\mathbb{R})$, $|C|$ denotes the 33-dimensional Lebesgue measure $\nu(C)$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$:
$f(x_{1},x_{2})=c\Big\{e^{-x_{1}^{2}-x_{2}^{2}}+\dfrac{1}{2}e^{-(x_{1}-2)^{2}-(x_{2}-2)^{2}}\Big\},$
where $c=\dfrac{2119}{9970}$. Let $b_{0}=\cfrac{1}{40}$, $b_{1}=\dfrac{1}{15}$, $b_{2}=c\big(\dfrac{1}{e^{8}}+\dfrac{1}{2}\big)$, $b_{3}=c\big(1+\dfrac{1}{2e^{8}}\big)$. We define $\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$ as
$\big[\underline{\kappa}_{1},\overline{\kappa}_{1}\big]:=\big[0,b_{0}\big],\quad\big[\underline{\kappa}_{2},\overline{\kappa}_{2}\big]:=\big[b_{0},b_{1}\big],\quad\big[\underline{\kappa}_{3},\overline{\kappa}_{3}\big]:=\big[b_{1},b_{2}\big],\quad\big[\underline{\kappa}_{4},\overline{\kappa}_{4}\big]:=\big[b_{1},b_{2}\big],\quad\big[\underline{\kappa}_{5},\overline{\kappa}_{5}\big]:=\big[b_{2},b_{3}\big].$
Moreover, we define $(c_{i,1},c_{i,2})$ and $d_{i}$ $(i=3,4,5)$ as
$(c_{3,1},c_{3,2})=(c_{5,1},c_{5,2})=(0,0),\quad(c_{4,1},c_{4,2})=(2,2),\quad d_{3}=\dfrac{5}{4},\quad d_{4}=d_{5}=1.$
Let $N=5$. The pattern blocks of $f(x_{1},x_{2})$ are defined as $B_{i}=E_{i}\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$, where
$E_{1}=E,\quad E_{2}=\left\{(x_{1},x_{2})\in E\ \middle|\ f(x_{1},x_{2})\geq b_{0}\right\},$
$E_{i}=\left\{(x_{1},x_{2})\in E\ \middle|\ (x_{1}-c_{i,1})^{2}+(x_{2}-c_{i,2})^{2}\leq d_{i}^{2}\right\}\qquad(i=3,4,5).$
Figure 5: $f(x_{1},x_{2})=b_{0}$
Figure 6: $f(x_{1},x_{2})=b_{1}$
Figure 7: $f(x_{1},x_{2})=b_{2}$
We assume that the independent random variables $U$, $\xi_{1}$, $\xi_{2}$, and $\eta$, defined on the probability space $(\Omega,\mathcal{F},P)$, are uniformly distributed on $[0,1)$. Let
$V_{1}^{(j)}=\underline{a}+(\overline{a}-\underline{a})\cdot\xi_{j}\quad(j=1,2),\quad W_{1}=b_{0}\cdot\eta.$ (18)
Subsequently, $(V_{1}^{(1)},V_{1}^{(2)},W_{1})$ satisfies (9). Let
$\widetilde{V}_{2}^{(1)}:=\dfrac{11}{2}\xi_{1}-2,\quad\widetilde{V}_{2}^{(2)}:=\dfrac{11}{2}\xi_{2}-2.$
We define $(V_{2}^{(1)},V_{2}^{(2)},W_{2})$ as
$(V_{2}^{(1)},V_{2}^{(2)}):=\left(\widetilde{V}_{2}^{(1)},\widetilde{V}_{2}^{(2)}\right)\big|_{E_{2}},\quad W_{2}:=(b_{1}-b_{0})\eta+b_{0}.$ (19)
Subsequently, $(V_{2}^{(1)},V_{2}^{(2)},W_{2})$ satisfies (9). Let
$\Phi_{i}(r,\theta,y):=(c_{i,1}+r\cos\theta,c_{i,2}+r\sin\theta,y),\quad\widetilde{B}_{i}:=\left[0,d_{i}\right]\times[0,2\pi)\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]\qquad(i=3,4,5).$
Thus, $B_{i}=\Phi_{i}(\widetilde{B}_{i})$ $(i=3,4,5)$. Let
$F_{R_{i}}(r):=\left\{\begin{array}[]{cl}0&(r\leq 0)\\ \dfrac{\pi r^{2}\big(\overline{\kappa}_{i}-\underline{\kappa}_{i}\big)}{|B_{i}|}&\left(0\leq r\leq d_{i}\right)\\ 1&\left(r\geq d_{i}\right).\end{array}\right$.
We define $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ as
$V_{i}^{(1)}=c_{i,1}+F_{R_{i}}^{-1}(\xi_{1})\cos 2\pi\xi_{2},\quad V_{i}^{(2)}=c_{i,2}+F_{R_{i}}^{-1}(\xi_{1})\sin 2\pi\xi_{2},\quad W_{i}=(\overline{\kappa}_{i}-\underline{\kappa}_{i})\eta+\underline{\kappa}_{i}\quad(i=3,4,5).$ (20)
Subsequently, $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ satisfy (9). Thus, by applying the pattern block method, we obtained $500000$ random vectors from the probability density function $f(x_{1},x_{2})$. To obtain $500000$ random vectors, $1375334$ random number quartets $(U,V_{\Psi(U)}^{(1)},V_{\Psi(U)}^{(2)},W_{\Psi(U)})$ were required. The adoption rate is calculated as follows:
$R_{\rm{adoption}}=\dfrac{1}{\sum_{i=1}^{5}|B_{i}|}=0.3644\cdots.$
Figs. 9 and 9 show the density scale histogram for $500000$ random vectors and the density function $y=f(x_{1},x_{2})$.
Figure 8: Histogram
Figure 9: $y=f(x_{1},x_{2})$
5 Conclusion
This study introduced the pattern block method, a generalization of the Ziggurat method. Future work will focus on developing an algorithm that identifies the pattern blocks of a given density function with high adoption rate and low search cost.
Acknowledgments
The authors thank Prof. Matsuura Yusuke (National Institute of Advanced Industrial Science and Technology), Prof. Masaaki Fukasawa (Osaka University) and Prof. Suguru Yamanaka (Aoyama Gakuin University) for their helpful comments and discussions on this study. We also thank Editage (www.editage.com) for English language editing. This work was supported by JSPS KAKENHI Grant Numbers JP22K01556 and JP25K05168.